1/ a/ The reason for split ratings across rating agencies and effects of split rating on...

1/

a/ The reason for split ratings across rating agencies and effects of split rating on financial market prices and agencies rating dynamics. Support your argument by example and evidence from recent literature.

b/ Explain the key uses of credit rating by regulators and financial market participants.

c/Explain, give example, the principle of cash matching technique for a bond portfolio.
Answer #1

Part (a)

Split ratings mean the same instruments has been rated differently by two different credit rating agencies. There can be multiple reasons for the same:

  1. The rating methodologies and principles followed by the two agencies may differ
  2. They may be using different rating scales
  3. They might have access to different set of information.
  4. They might have interpreted subjective information differently.
  5. They might have different views on the sector, industry and the company.
  6. They might have taken different assumptions pertaining to the business.
  7. They opinions might have different biases within themselves

Effects of split rating on financial market prices and agencies rating dynamics:

  1. Preference to agencies: Companies and organizations will be attracted towards the agencies that rate them better. This will lead to loss of neutrality in the business of credit of rating. This will give rise to unwanted practices in the rating business.
  2. Ratings play a significant role in investment decisions. The investment decisions may not yield the expected results.
  3. Ratings play a significant role in calculation of the safety capital by the banks, split ratings can lead to over / understatement in the calculation of safety capital by the banks.
  4. Confusion amongst the investor community: Different ratings of the same instrument will lead to enormous chaos and disorder amongst the investment community.

There is no shortage of literature on the topics of split ratings. One example of importance will be to examine the considerations for split rating given under Basel II.

Under Basel II:

  1. Banks are not allowed to pick and choose their rating agencies.
  2. Banks are required to be consistent in the choice and usage of rating agencies.
  3. Banks are required to choose a consistent rating policy
  4. In fact, the need to induce refinements in the existing regulations was felt due to presence of split ratings and adverse impacts of split ratings on financial market prices and agencies rating dynamics.

Part (b)

Explain the key uses of credit rating by regulators and financial market participants

  • Credit rating is an independent assessment of creditworthiness of the issuer of the bond, be it a company, organisation, special purpose vehicle or even government. Financial market participants (traders, investors, lenders etc) use this rating to make their own investment decisions.
  • Financial market participants make use of the credit ratings to understand the risk associated with the investment security. This in turn helps them decide on the expected return from the security.
  • Regulators use the credit ratings to assess the riskiness of the new securities being issued, to ensure safety from the perspective of retail investors. Thus regulators insist on rating of all IPOs, bond issuance, new loans, new mortgage backed securities.
  • Issuers of instruments use credit ratings to improve the marketability of the securities being issued.

Part (c)

Cash (flow) matching technique is an investment strategy where an investor invests in securities that generate stream of cash flows that matches the liabilities or cash flow requirements of the investor. This is another way of saying that i choose my investments in such a way that they mature and return me the money exactly as and when i require them. Thus, this exploits the principal of matching the maturity of investments to that of liabilities.

An example:

Let's say I require $ 1,100,000 at the end of year 1 and $ 640,000 at the end of year 2. Let's further assume that the 3 year bond pays an annual coupon of 10% while a 2 year maturity bond pays an annual coupon of 8%. What should be my investment strategy?

Solution:

At the end of year 3, the 3 year maturity bond will return Principal as well as interest = $ 1000 + 10% x 1000 = $ 1,100

So, I need to invest in 1,100,000 / 1,100 = 1000 nos. of 3 year maturity bond paying annual coupon of 10%.

These bonds will also pay a total coupon of 1,000 x 1,000 x 10% = $ 100,000 at the end of year 2. My total requirement at the end of year 2 = $ 640,000

Of which, amount matched by coupon from 3 year maturity bond = $ 100,000

Balance requirement = 640,000 - 100,000 = $ 540,000

Proceeds from one 2 year maturity bond at the of year 2 = 1000 x (1 + 8%) = $ 1,080

Nos. of 2 year maturity bonds required = 540,000 / 1,080 = 500

Thus i need to invest in 500 numbers of 2 year maturity bonds and 1000 numbers of 3 year maturity bonds to match my cash flow requirement.